OpenStudy (anonymous):

If sinθ =2/3 , which of the following are possible? A.secθ=-3/2 and tanθ=2/√5 B.cosθ=√5/3 and tanθ= 2/√5 C.secθ= 3/√5 and tanθ= 2/√5 D.cosθ=-√5/3 and tanθ=2/√5

OpenStudy (anonymous):

2,3 sqrt 5 triangle.....

OpenStudy (anonymous):

C.secθ= 3/√5 and tanθ= 2/√5

OpenStudy (anonymous):

and B.cosθ=√5/3 and tanθ= 2/√5

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

\[\sin(\theta)=\frac{2}{3}\] \[\sin^2(\theta)=(\frac{2}{3})^2\] \[\sin^2(\theta)=\frac{4}{9}\] ------------------------------------------------ \[\sin^2(\theta)+\cos^2(\theta)=1\] \[\frac{4}{9}+\cos^2(\theta)=1\] \[\cos^2(\theta)=1-\frac{4}{9}\] \[\cos^2(\theta)=\frac{5}{9}\] \[\cos(\theta)=\sqrt{\frac{5}{9}}\] or \[\cos(\theta)=-\sqrt{\frac{5}{9}}\] \[\cos(\theta)=\frac{\sqrt{5}}{3}\] or \[\cos(\theta)=-\frac{\sqrt{5}}{3}\] \[\sec(\theta)=\frac{3}{\sqrt{5}}\] or \[\sec(\theta)=-\frac{3}{\sqrt{5}}\] ----------------------------------------------- \[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\] \[\tan(\theta)=\frac{\frac{2}{3}}{\frac{\sqrt{5}}{3}}\] \[\tan(\theta)=\frac{2}{\sqrt{5}}\] So the answer is choice C

OpenStudy (anonymous):

tsk, tsk, gotta watch those pesky quadrants....